Given z=cos A + i sin A and u+iv =(1+z)(1+z^2).Prove that v=u tan (3A/2),u^2+ v^2=16 cos^2(A/2)cos^2(A).Hint:A=theta
1
Expert's answer
2020-05-05T18:36:29-0400

u+iv = (1+z)(1+z2) = 1 +z2 +z + z3

Since we know that if z = cosA = i sinA, then, zn = cos(nA) + i sin(nA)

u+iv = 1 + cos2A + cosA + cos3A + i( sin2A + sinA + sin3A)

Comparing both sides, we get,

u= 1 + cos2A + cosA + cos3A

v= sin2A + sinA + sin3A


To prove: v = u tan(3A/2), it is enough to prove u. sin (3A/2) = v. cos (3A/2)


u. sin (3A/2)= (1 + cos2A + cosA + cos3A)sin (3A/2)

= sin (3A/2) + cos2A. sin (3A/2) + cosA. sin (3A/2) + cos3A. sin (3A/2) equation 1

Using identity 2.sin(x) . cos(y) = sin (x+y) + sin (x-y)

= (1/2) sin (3A/2) + (1/2) sin (7A/2) + (1/2) sin (5A/2) + (1/2) sin (9A/2)


v. cos (3A/2) = (sin2A + sinA + sin3A) . cos (3A/2)

= sin2A . cos (3A/2)+ sinA. cos (3A/2) + sin3A. cos (3A/2)

Using identity 2.sin(x) . cos(y) = sin (x+y) + sin (x-y)

= (1/2) sin (3A/2) + (1/2) sin (7A/2) + (1/2) sin (5A/2) + (1/2) sin (9A/2) equation 2

From equation1 and equation2,

u. sin (3A/2) = v. cos (3A/2)

v = u tan(3A/2), hence, proved.



To prove: u2 + v2 = 16 cos2(A/2) cos2(A/2)

u2 = (1 + cos2A + cosA + cos3A)2

Using identities (a+b+c+d)2 = a2 +b2 + c2 +d2 + 2(ab+bc+cd+da+ac+bd) and

2.cos(x) . cos(y) = cos (x+y) + cos (x-y)

=1 + cos2(2A) + cos2A + cos2(3A) + 3 cos (2A) + 3 cos(3A) + cos(5A) + 4 cos(A) equation3


v2 = ( sin2A + sinA + sin3A)2

Using identities (a+b+c)2 = a2 +b2 + c2 + 2(ab+bc+ac) and

2.sin(x) . sin(y) = cos (x-y) - cos (x+y)

= sin2 (2A) + sin2(A) + sin2(3A) + 2cosA - cos(3A) + cos(2A) - cos(4A) - cos(5A) equation4


Using identity sin2(x) + cos2(x) = 1,from equations 3 and 4 ,

u2 + v2 = 4+4cos(2A)+2cos(3A)+6cos(A)

= 4(1+cos(2A)) + 4cos(A)(1+cos(2A))

= 4 (1+cos(A)) . (1+cos(2A))

Using 2cos2(x) = 1+ cos(2x)

= 16 cos2(A) cos2(A/2)